OpenStudy (anonymous):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle? My answer is: 3. Use the Pythagorean Theorem to find the lengths of the triangle and apply those answers to the Law of Sines. Two lengths need to be known. i need help creating an example. @tomo

OpenStudy (anonymous):

@Mertsj

jimthompson5910 (jim_thompson5910):

you could also use the trig functions sine, cosine, tangent etc

OpenStudy (anonymous):

I think what theyre looking for is like give an example using a right trianglee

jimthompson5910 (jim_thompson5910):

for example, if you know the opposite and adjacent legs (and let's say they are 3 and 4), then you can say tan(x) = opp/adj tan(x) = 3/4 x = arctan(3/4) x = 36.8698976 degrees

OpenStudy (anonymous):

 that right?

jimthompson5910 (jim_thompson5910):

yes, so if you wanted to know this angle here  then you would solve tan(x) = 3/4 for x

jimthompson5910 (jim_thompson5910):

x is not a side length in my pic , it's an angle

OpenStudy (anonymous):

makes sense :)

jimthompson5910 (jim_thompson5910):

that's great, glad it does